# Unified Multimodal Understanding and Generation Models: Advances, Challenges, and Opportunities

### v3

@@ 4.2 Text-to-Image Datasets
  • Synthesized Datasets: Specialized datasets for text-to-image synthesis are increasingly created using existing generative models. The text-to-image-2M dataset [228] provides  2 million enhanced text-image pairs for fine-tuning, curated using advanced T2I and captioning models. SFHQ-T2I [230] offers 122K diverse, high-resolution synthetic face images generated by multiple T2I models, ensuring variance and privacy. For entity control, the EliGen TrainSet [231] uses images from a baseline model (FLUX.1-dev) and MLLM-generated prompts for stylistic consistency and detailed annotation. Lastly, BLIP-3o 60k [159] provides 60,000 instruction tuning samples distilled from GPT-4o, covering various categories for diverse training.
@@ 4.5 Other Text+Image-to-Image Datasets
MagicBrush [234] $>$10K Edits Real‑image Editing Image + Instruction $\rightarrow$ Image NeurIPS2023
EditVal [294] N/A Attribute‑focused Eval Image + Instruction $\rightarrow$ Image Arxiv2023
Emu-Edit [295] 5.5K Edits Multi‑task Editing Image + Instruction $\rightarrow$ Image CVPR2024
@@
I2EBench [297] $>$4K Edits Multi‑dimensional Eval Image + Instruction $\rightarrow$ Image NeurIPS2024
@@ 5.2 Evaluation on Image Generation
To support large-scale concept diversity, EvalMuse-40K [316] provides 40K crowdsourced prompts focusing on nuanced concept representation, and HEIM [286] identifies 12 aspects, including text-image alignment, image quality, aesthetics, originality, reasoning, knowledge, bias, toxicity, fairness, robustness, multilinguality and efficiency. Considering practical needs, FlashEval [317] shrinks the large-scale evaluation set into diverse smaller ones through iterative search to accelerate the benchmark testing. MEMO-Bench [318] introduces a comprehensive benchmark for evaluating the emotional understanding and expression capabilities of T2I models and MLLMs. ConceptMix [290] evaluates text-to-image models’ compositional generation ability by sampling k-tuples of visual concepts to construct prompts and automatically verifying concept presence in the resulting images using a strong visual language model. TIFA [283] offers a fine-grained benchmark for evaluating text-to-image faithfulness via visual question answering generated from prompts. To enrich dependencies in question generation for VQA-based evaluation of image–prompt alignment, DSG-1k [288] refines its questions using a multi-level semantic graph.
@@
More recent benchmarks include GEdit-Bench [300] which features 606 real-world instruction–image pairs, CompBench [301] that decomposes edits into location, appearance, dynamics and object dimensions via large-scale MLLM–and–human collaboration, and GIE-Bench [302] which uses multiple-choice VQA and object-aware masking on over 1,000 examples to evaluate editing accuracy and content preservation.

### v4

@@ 4.1 Multimodal Understanding Datasets
  • ALLaVA [209]: This dataset, comprising 1.4 million samples, is synthetically generated to facilitate the training of resource-friendly Lite Vision-Language Models (LVLMs). The generation pipeline leverages strong proprietary models (like GPT-4V) in a multi-stage process: first, images are selected from sources like LAION and Vision-FLAN; then, fine-grained, detailed captions are generated for these images; finally, complex reasoning visual question-answering pairs are created, emphasizing detailed answers that include evidence and chain-of-thought, to support robust visual instruction fine-tuning.
@@ 4.2 Text-to-Image Datasets
  • DOCCI [250]: DOCCI provides 15k uniquely curated images, each with long, human-annotated English descriptions (average 136 words) designed to be highly detailed and to differentiate between similar images. The dataset’s focus on fine-grained descriptions and contrastive image sets makes it a valuable resource for training and evaluating both image-to-text and text-to-image models, particularly for their ability to handle nuanced details and complex compositions.
@@
  • Synthesized Datasets: Specialized datasets for text-to-image synthesis are increasingly created using existing generative models. The text-to-image-2M dataset [254] provides  2 million enhanced text-image pairs for fine-tuning, curated using advanced T2I and captioning models. SFHQ-T2I [256] offers 122K diverse, high-resolution synthetic face images generated by multiple T2I models, ensuring variance and privacy. For entity control, the EliGen TrainSet [257] uses images from a baseline model (FLUX.1-dev) and MLLM-generated prompts for stylistic consistency and detailed annotation. Similarly, BLIP-3o 60k [165] provides 60,000 instruction tuning samples distilled from GPT-4o, covering various categories for diverse training. Lastly, ShareGPT-4o-Image[259] contributes 45K text-to-image pairs, where prompts are generated through both a structured attribute-first approach and an image-first approach, with corresponding images synthesized by GPT-4o’s image generation capabilities to distill its advanced skills.
@@ 4.3 Image Editing Datasets
  • HIVE [262]: The HIVE framework introduces human feedback into instructional visual editing, providing a 1.1M training dataset (generated similarly to InstructPix2Pix using GPT-3 and Prompt-to-Prompt, plus cycle consistency augmentation) and a 3.6K reward dataset where humans rank model outputs.
  • EditWorld [264]: EditWorld introduces the task of “world-instructed image editing,” focusing on realistic world dynamics. Its dataset is curated through two branches: one uses GPT-3.5 for world instructions and T2I models for complex input-output image generation, and the other extracts paired frames from videos with vision-language models generating corresponding instructions for dynamic transformations.
  • PromptFix [266]: PromptFix constructs a large-scale instruction-following dataset ( 1.01M triplets) focusing on a comprehensive range of image processing tasks, particularly low-level tasks (e.g., inpainting, dehazing, super-resolution, colorization).
@@
  • RefEdit [269]: This synthetic dataset specifically targets instruction-based editing challenges involving referring expressions in complex scenes. It’s generated using GPT-4o for text components (prompts, instructions, referring expressions), FLUX for initial images, Grounded SAM for precise mask generation from expressions, and specialized models for controlled edits like object removal or modification.
  • ImgEdit [270]: ImgEdit is a large-scale (1.2M edit pairs) dataset designed for high-quality single-turn and multi-turn image editing. Its multi-stage generation pipeline filters LAION-Aesthetics, uses vision-language models and detection/segmentation models for grounding and instruction generation (including spatial cues and multi-turn dialogues), employs state-of-the-art generative models (FLUX, SDXL with plugins) for task-specific inpainting, and uses GPT-4o for final quality filtering.
  • ByteMorph-6M [271]: ByteMorph-6M is a large-scale dataset with over 6 million image editing pairs specifically designed for instruction-guided editing involving non-rigid motions (e.g., camera viewpoint shifts, object deformations, human articulations). It is constructed by first using a Vision-Language Model to generate “Motion Captions” from instruction templates for an initial frame; then, an image-to-video model generates a dynamic video based on this motion caption; finally, frames are sampled from these videos, and an LLM generates precise editing instructions describing the transformation between neighboring frame pairs, which form the source-target editing data.
  • ShareGPT-4o-Image (Editing) [259]: Complementing its text-to-image data, ShareGPT-4o-Image also includes 46K instruction-guided image editing triplets. These samples are generated by first selecting a source image (either from its text-to-image collection or real photos), then sampling an editing task from a predefined taxonomy (e.g., object manipulation, style transfer), having an LLM synthesize a natural language instruction for that task and image, and finally using GPT-4o’s image generation capabilities to produce the edited image.
@@ 4.4 Interleaved Image-Text Datasets
  • OmniCorpus [275]: OmniCorpus is a very large-scale (10 billion-level) image-text interleaved dataset, containing 8.6 billion images and 1,696 billion text tokens from 2.2 billion documents. It was created using an efficient data engine that extracts and filters content from diverse sources, including English and non-English websites as well as video platforms (extracting keyframes and transcribing audio). The dataset incorporates human-feedback filtering to enhance data quality, aiming to provide a solid foundation for MLLM research.
@@ 4.5 Other Text+Image-to-Image Datasets
  • Graph200K [280]: Graph200K is a graph-structured dataset built upon Subjects200K, where each image is augmented with 49 types of annotations spanning five meta-tasks: conditional generation (e.g., canny edges, depth maps, segmentation), IP preservation, style transfer (semantic-variant and -invariant), image editing (background-variant and -invariant using VLMs and inpainting models), and restoration (via online degradations). This structure aims to increase task density and inter-relation, enabling models to learn shared and transferable knowledge for universal image generation by formulating tasks as paths within this graph.
@@
MMIG-Bench [328] 4850 Prompts Comprehensive Eval Text Prompt $\rightarrow$ Image Arxiv2025
@@
MagicBrush [261] 1053 Edits Real‑image Editing Image + Instruction $\rightarrow$ Image NeurIPS2023
EditVal [330] 648 Edits Attribute‑focused Eval Image + Instruction $\rightarrow$ Image Arxiv2023
Emu-Edit [331] 3055 Edits Multi‑task Editing Image + Instruction $\rightarrow$ Image CVPR2024
@@
I2EBench [333] 2240 Edits Multi‑dimensional Eval Image + Instruction $\rightarrow$ Image NeurIPS2024
@@
AnyEdit [268] 1250 Edits Comprehensive Eval Image + Instruction $\rightarrow$ Image CVPR2025
@@
EditInspector [339] 783 Edits Comprehensive Eval Image + Instruction $\rightarrow$ Image Arxiv2025
ComplexBench-Edit [340] $<$1K List of Edits Chain-dependent Editing Eval Image + Instruction $\rightarrow$ Image Arxiv2025
ByteMorph-Bench [271] 613 Edits Non-rigid Editing Eval Image + Instruction $\rightarrow$ Image Arxiv2025
RefEdit-Bench [269] 200 Edits Expression-driven Editing Eval Image + Instruction $\rightarrow$ Image Arxiv2025
ImgEdit-Bench [270] 200 Edits Expression-driven Editing Eval Image + Instruction $\rightarrow$ Image Arxiv2025
KRIS-Bench [341] 1267 Edits Cognitive Reasoning Eval Image + Instruction $\rightarrow$ Image Arxiv2025
@@ 5.2 Evaluation on Image Generation
To support large-scale concept diversity, EvalMuse-40K [355] provides 40K crowdsourced prompts focusing on nuanced concept representation, and HEIM [321] identifies 12 aspects, including text-image alignment, image quality, aesthetics, originality, reasoning, knowledge, bias, toxicity, fairness, robustness, multilinguality and efficiency. Considering practical needs, FlashEval [356] shrinks the large-scale evaluation set into diverse smaller ones through iterative search to accelerate the benchmark testing. MEMO-Bench [357] introduces a comprehensive benchmark for evaluating the emotional understanding and expression capabilities of T2I models and MLLMs. ConceptMix [325] evaluates text-to-image models’ compositional generation ability by sampling k-tuples of visual concepts to construct prompts and automatically verifying concept presence in the resulting images using a strong visual language model. TIFA [318] offers a fine-grained benchmark for evaluating text-to-image faithfulness via visual question answering generated from prompts. To enrich dependencies in question generation for VQA-based evaluation of image–prompt alignment, DSG-1k [323] refines its questions using a multi-level semantic graph. MMIG-Bench [328] introduced a multi-dimensional assessment framework that rigorously examines text-to-image generation models.
@@
More recent benchmarks include GEdit-Bench [336] which features 606 real-world instruction–image pairs, CompBench [337] that decomposes edits into location, appearance, dynamics and object dimensions via large-scale MLLM–and–human collaboration, and GIE-Bench [338] which uses multiple-choice VQA and object-aware masking on over 1,000 examples to evaluate editing accuracy and content preservation. Following this trend, benchmarks like [340, 339, 268, 270] also undertake comprehensive evaluation of text-guided image editing, assessing vision consistency, artifact detection, instruction adherence, visual quality, and detail preservation.
Other benchmarks include ByteMorph-Bench [271] which tackles non-rigid image manipulation, RefEdit-Bench [269] which evaluates referring-expression–based edits in complex multi-entity scenes, and KRIS-Bench [341] which offers a cognitively grounded suite assessing factual, conceptual and procedural reasoning.